Sodium atoms emit a spectral line with a wavelength in the yellow, 589.6mm. What is the difference in energy between the two energy levels involved in the emission of this spectral line? Answer: 2.1eV


1
Expert's answer
2021-03-12T09:22:46-0500

We know that the energy of emitted photon is equal to the difference of the energies of initial and final energy levels. The energy of photon is

"E = h\nu = h\cdot \dfrac{c}{\lambda} = 6.63\cdot10^{-34}\,\mathrm{J\cdot s}\cdot \dfrac{3\cdot10^8\,\mathrm{m/s}}{589.6\cdot10^{-6}\,\mathrm{m}} = 3.4\cdot 10^{-19}\,\mathrm{J} \approx 2.1\,\mathrm{eV}."